Calculate the following in 5.28 g Li3PO4 a) Li ions b) grams O
Answer #1

Molar mass of Li3PO4 = 115.79 g/mol

Atomic weight of Li = 6.94 g/mol

Atomic weight of O = 16 g/mol

Number of moles = given mass in gram / Molar mass

Number of moles of Li3PO4 = 5.28 g / 115.79 g/mol = 0.0456 mol

Since number of Li ions in Li3PO4 = 3 so number of moles of Li ions = 3 * 0.0456 mol = 0.1368 mol

Since number of O in Li3PO4 = 4 hence number of moles of O = 4 * 0.0456 mol = 0.1824 mol

a) Mass of Li ions = moles of Li ions * Atomic weight of Li = 0.1368 mol * 6.94 g/mol = 0.95 g

b) Mass of O = moles of O * Atomic weight of O = 0.1824 mol * 16 g/mol = 2.92 g